40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 29 11:51:27 2021, Anchal, Summary, LSC, Start of measuring IMC WFS noise contribution in ar cavity length noise 
    Reply  Thu Apr 29 15:11:33 2021, gautam, Update, CDS, RFM  RFM_errs.pngScreenshot_2021-04-29_15-12-56.png
       Reply  Thu Apr 29 17:43:16 2021, Koji, Update, CDS, RFM  
       Reply  Thu Apr 29 17:43:48 2021, Anchal, Update, CDS, F2A Filters double check F2AFiltersON.png
          Reply  Fri Apr 30 00:20:30 2021, gautam, Update, CDS, F2A Filters double check 
    Reply  Thu Apr 29 17:51:19 2021, Anchal, Summary, LSC, Start of measuring IMC WFS noise contribution in arm cavity length noise WFS1_PIT_exc_80-100Hz_Arms_ASD.pdf
       Reply  Fri Apr 30 00:18:40 2021, gautam, Summary, LSC, Start of measuring IMC WFS noise contribution in arm cavity length noise 
       Reply  Mon May 3 09:14:01 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise WFS1_PIT_Noise_Inj_Test_IMC_unlocked.pdfWFS1_PIT_Noise_Inj_Test_ARM_locked.pdf
          Reply  Mon May 3 17:28:58 2021, Anchal, Paco, Rana, Update, LSC, IMC WFS noise contribution in arm cavity length noise excitationoftheMCanglessothatwecanseesomethingdotpng.pngexcitationoftheMCanglessothatwecanseesomethingdotpngbutthistimeitsMC2.png
             Reply  Tue May 4 11:43:09 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise ArmCavNoiseContributions.pdfIOO-MC1_PIT_NoiseInjTest2.pdfIOO-MC1_PIT_NoiseInjTest_AWGGUI_Config.png
                Reply  Fri May 7 11:54:02 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise ITMX-XARM_TF.pdfITMY-YARM_TF.pdfArmCavNoiseContributions.pdf
                   Reply  Mon May 10 18:19:12 2021, Anchal, Paco, Update, LSC, IMC WFS noise contribution in arm cavity length noise, Corrections ArmCavNoiseContributions.pdf
                      Reply  Wed May 12 10:53:20 2021, Anchal, Paco, Update, LSC, PSL-IMC PDH Loop and XARM PDH Loop diagram IMC_SingleArm.pdf
Message ID: 16095     Entry time: Thu Apr 29 11:51:27 2021     Reply to this: 16097   16101
Author: Anchal 
Type: Summary 
Category: LSC 
Subject: Start of measuring IMC WFS noise contribution in ar cavity length noise 

Tried locking the arms

  • Ran IFO > Configure > ! (YARM) > Restore YARM. Nothing happened.
  • Trying to align through tip-tilt:
    • Previous values: TT1: PIT: -1.7845, YAW: -0.2775; TT2: PIT: -1.3376, YAW: -0.0436
    • Couldn't get flashing of light in the arms at all.
    • Restored values to previous values.
  • Noticed that ITMY OPLEV YAWW Error has gone very high overnight while other oplevs remained the same.
  • Trying to change the C1:SUS-ITMY_YAW_OFFSET to bring this oplev yaw error back to near zero.
  • Changed C1:SUS-ITMY_YAW_OFFSET from -34 to 50. OPLEV_YEROR reduced to around -23 from -70.
  • Same thing with BS PIT. OPLEV_PERROR is highlighted in red at -52.
  • Changing C1:SUS-BS_PIT_OFFSET from 55 to 30. This brought OPLEV_PERROR to -15 from -52.
  • Trying to align PRM by changing C1:SUS-PRM_PIT_OFFSET and C1:SUS-PRM_YAW_OFFSET.
  • Inital values: C1:SUS-PRM_PIT_OFFSET: -20 , C1:SUS-PRM_YAW_OFFSET: 39.

Did the WFS step response test on IMC in between while waiting for help. See 16094.


Back to trying arm locking

  • Tried IFO > Configure > ! (XARM) > Restore YARM. Nothing happened.
  • Tried IFO > Configure > ! (YARM) > Restore YARM. Nothing happened again.
  • Tried Movie Capture of AS screen from VIDEO > Movie Capture > AS. But the script failed due to module not present error.

PMC got unlocked

  • Infront of me, PMC got unlocked. I did not go to PMC locking the screen at all since morning.
  • I opened the C1PSL_PMC screen. The PSL Autolocker blinker is not blinking but the switch is set to Enable. 
  • I do not see any automatic effort happening for regaining lock at PMC.
  • I'll try it manually. I was able to get the PMC locked again. C1:PSL-PMC_PMCTRANSPD is showing 0.761 V and C1:PSL-PMC_RFPDDC is showing 0.053 V.
  • Now IMC auto locker seems to be trying to get the lock acquired.
  • It acquired a lock a few times but struggled to keep it on. I reduced C1:IOO-WFS_GAIN to 0 and then the lock could stay on. Seemed like the accumulated offsets were not good.
  • So I cleared the history on WFS1, TRANS and WFS2 filter banks and then ramped the WFS overall gain (C1:IOO-WFS_GAIN) back to 1 and now IMC seems to stay locked in a stable configuration.
  • However, I still don't know what caused the PMC to get unlocked in the first place. Did my repeated arm locking attempts did something to the main laser frequency?

Back to trying arm locking

  • Tried IFO > Configure > ! (YARM) > Restore YARM again. Nothing happened again.
ELOG V3.1.3-